# Classifying representations of finite classical groups of Lie type of dimension up to $\ell^{4}$

### v1

@@ 2.3. Multiplicity and dimension
following [8, §10.2 and §13.3].
Now, let $V(\lambda)$ be a finite-dimensional $G(\overline{\mathbb{F}}_{p})$-module with highest weight $\lambda$ and $\mu\preceq\lambda$ be a dominant weight. We define the multiplicity of $\mu$ in $V(\lambda)$ as
$m_{\lambda}(\mu)=\dim V(\lambda)_{\mu}$
which can be calculated by using the recursive Freudenthal multiplicity formula [8, §22.3]:
$m_{\lambda}(\mu)=\frac{2\sum_{\alpha\in\Phi_{+}}\sum_{i=1}^{\infty}(\mu+i\alpha,\alpha)m_{\lambda}(\mu+i\alpha)}{\parallel\lambda+\delta\parallel^{2}+\parallel\mu+\delta\parallel^{2}}=\frac{2\sum_{\alpha\in\Phi_{+}}\sum_{i=1}^{\infty}(\mu+i\alpha,\alpha)m_{\lambda}(\mu+i\alpha)}{(\lambda,\lambda)-(\mu,\mu)+2(\lambda-\mu,\delta)}$
@@
(2.1) $\dim V(\lambda)=\sum_{\mu\in\Pi(\lambda)}m_{\lambda}(\mu)l(\mu)$
3. Multiplicities and dimensions for groups of type $A_{\ell}$
In order to be ready for computing certain multiplicities required along the section, let us start by giving a system of simple roots of type $A_{\ell}$ and the fundamental weights associated.
Let $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{\ell}\}$ be the simple roots system of type $A_{\ell}$ given by
@@ 3. Multiplicities and dimensions for groups of type $A_{\ell}$
$\lambda_{i}=\frac{1}{\ell+1}[(\ell-i+1)\alpha_{1}+2(\ell-i+1)\alpha_{2}+i(\ell-i+1)\alpha_{i}+i(\ell-i)\alpha_{i+1}+\cdots+i\alpha_{\ell}].$
The following result will useful for computing the dimensions at hand.
@@
For any $1\leq i,j\leq\ell$ we have that
$(\ell+1)^{2}(\lambda_{i},\lambda_{j})$ $=i(j-1)(\ell-j+1)^{2}\cdot(-1/2)+ij\sum_{k=1}^{\ell-j+1}k.$
@@
Without loss of generality suppose that $i\leq j$, namely $j=i+r$ with $r\geq 0$. Thus, we have
$\lambda_{i}$ $=(\ell-i+1)\alpha_{1}+\cdots+i(\ell-i+1)\alpha_{i}+\cdots+i(\ell-i-r+1)\alpha_{j}+i(\ell-i-r)\alpha_{j+1}+\cdots+i\alpha_{\ell}$
$\lambda_{j}$ $=(\ell-j+1)\alpha_{1}+\cdots+(\ell-j+1)\alpha_{i}+\cdots+j(\ell-j+1)\alpha_{j}+j(\ell-j)\alpha_{j+1}+\cdots+j\alpha_{j}.$
Then we can calculate
$(\ell+1)^{2}(\lambda_{i},\lambda_{j})$
$=(\ell-i+1)(\ell-j+1)+2(\ell-i+1)(\ell-j+1)\cdot(-1/2)$
$+2(\ell-i+1)(\ell-j+1)\cdot(-1/2)+4(\ell-i+1)(\ell-j+1)+6(\ell-i+1)(\ell-j+1)\cdot(-1/2)$
$+6(\ell-j+1)(\ell-i+1)\cdot(-1/2)+9(\ell-i+1)(\ell-j+1)+12(\ell-j+1)(\ell-i+1)\cdot(-1/2)$
$\vdots$
$+i(j-2)(\ell-j+1)(\ell-i-r+2)\cdot(-1/2)+i(j-1)(\ell-i-r+2)(\ell-j+1)+ij(\ell-j+1)(\ell-i-r+2)\cdot(-1/2)$
$+i(j-1)(\ell-i-r+1)(\ell-j+1)\cdot(-1/2)+ij(\ell-i-r+1)(\ell-j+1)+ij(\ell-i-r+1)(\ell-j)\cdot(-1/2)$
$+ij(\ell-i-r)(\ell-j+1)\cdot(-1/2)+ij(\ell-i-r)(\ell-j)+ij(\ell-i-r)(\ell-j-1)\cdot(-1/2)$
$+ij(\ell-i-r-1)(\ell-j)\cdot(-1/2)+ij(\ell-j-1)^{2}+ij(\ell-i-r-1)(\ell-j-2)\cdot(-1/2)$
$\vdots$
$+6ij\cdot(-1/2)+4ij+2ij\cdot(-1/2)$
$+2ij\cdot(-1/2)+ij$
$=i(j-1)(\ell-j+1)^{2}\cdot(-1/2)+ij(\ell-j+1)^{2}+ij(\ell-j+1)(\ell-j)\cdot(-1/2)$
$+ij(\ell-j)(\ell-j+1)\cdot(-1/2)+ij(\ell-j)^{2}+ij(\ell-j)(\ell-j-1)\cdot(-1/2)$
$+ij(\ell-j-1)(\ell-j)\cdot(-1/2)+ij(\ell-j-1)^{2}+ij(\ell-j-1)(\ell-j-2)\cdot(-1/2)$
$\vdots$
$+6ij\cdot(-1/2)+4ij+2ij\cdot(-1/2)$
$+2ij\cdot(-1/2)+ij$
$=i(j-1)(\ell-j+1)^{2}\cdot(-1/2)+ij\sum_{k=1}^{\ell-j+1}k^{2}-\sum_{k=1}^{\ell-j+1}k(k-1)$
$=i(j-1)(\ell-j+1)^{2}\cdot(-1/2)+ij\sum_{k=1}^{\ell-j+1}[k^{2}-k(k-1)]$
$=i(j-1)(\ell-j+1)^{2}\cdot(-1/2)+ij\sum_{k=1}^{\ell-j+1}k.$
Now we introduce a pivotal concept, specially in the next two sections.
Definition 3.2.
Let $G$ be a connected reductive simple algebraic group of type $A_{\ell}$ over $\overline{\mathbb{F}}_{p}$ of simply connected type and $\lambda$ be a dominant weight. We will say that $\lambda$ is admissible if the dimension of the $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ is at most $\ell^{4}/16$.
We see for instance.
@@
When $G$ is as above with $\ell>2$, the dominant weight $\lambda_{1}+\lambda_{2}$ is admissible.
We need to compute the dimension of $V(\lambda_{1}+\lambda_{2})$ to verify the admissibility of $\lambda=\lambda_{1}+\lambda_{2}$. Observe first that $\lambda_{3}=\lambda-(\alpha_{1}+\alpha_{2})$, we then have to compute $m_{\lambda}(\lambda_{3}+\alpha_{1})$ and is idem for $m_{\lambda}(\lambda_{3}+\alpha_{2})$; we set $\mu=\lambda_{3}+\alpha_{2}=\lambda-\alpha_{1}$. Thus we have that
$m_{\lambda}(\mu)=2\cdot\frac{m_{\lambda}(\mu+\alpha_{2})(\mu+\alpha_{2},\alpha_{2})}{(\lambda+\delta,\lambda+\delta)-(\mu+\delta,\mu+\delta)}.$
Immediately we see that $m_{\lambda}(\mu+\alpha_{2})=m_{\lambda}(\lambda)=1$ and $(\mu+\alpha_{2},\alpha_{2})=1/2$. Thanks to Lemma 3.1 we deduce that
$(\lambda,\lambda)=\frac{5\ell-4}{2(\ell+1)}\quad\text{ and }\quad(\mu,\mu)=\frac{5\ell-4}{2(\ell+1)}.$
We then have
$m_{\lambda}(\mu_{2})$ $=2\cdot\frac{m_{\lambda}(\mu+\alpha_{2})(\mu+\alpha_{2},\alpha_{2})}{(\lambda,\lambda)-(\mu,\mu)+2(\lambda-\mu,\delta)}$
$=2\cdot\frac{1\cdot\frac{1}{2}}{\frac{5\ell-4}{2(\ell+1)}-\frac{5\ell-4}{2(\ell+1)}+1}$
$=1.$
Hence we compute the multiplicity of $\mu$ in $V(\lambda)$ as follows
$m_{\lambda}(\mu)$ $=\frac{(\mu+\alpha_{1},\alpha_{1})m_{\lambda}(\mu+\alpha_{1})+(\mu+\alpha_{2},\alpha_{2})m_{\lambda}(\mu+\alpha_{2})+(\mu+\alpha_{1}+\alpha_{2},\alpha_{1}+\alpha_{2})m_{\lambda}(\mu+\alpha_{1}+\alpha_{2})}{(\lambda+\delta,\lambda+\delta)-(\mu+\delta,\mu+\delta)}$
$=2\cdot\frac{1\cdot 1+1\cdot 1+1\cdot 1}{\frac{5\ell-4}{2(\ell+1)}-\frac{3(\ell-2)}{2(\ell+1)}+2}$
$=2.$
Thus, we conclude that
$\dim V(\lambda_{1}+\lambda_{2})=m_{\lambda}(\lambda)l(\lambda)+m_{\lambda}(\lambda_{3})l(\lambda_{3})$ $=\ell(\ell+1)+2\cdot\frac{(\ell-1)\ell(\ell+1)}{6}$
$=\ell(\ell+1)[1+(\ell-1)/3]$
$=\ell(\ell+1)(\ell+2)/3,$
as desired.
In contrast with example above we observe.
Example 3.4.
For all $\ell\geq 12$, the weight $\lambda=\lambda_{\ell-2}+\lambda_{\ell}$ is not admissible.
It suffices to show that the multiplicity of $\lambda_{\ell-3}\preceq\lambda$ is $3$ by resorting as follows
$\dim V(\lambda_{\ell-2}+\lambda_{\ell})\geq m_{\lambda}(\lambda_{\ell-3})l(\lambda_{\ell-3})$ $=3\cdot\frac{(\ell-2)(\ell-1)\ell(\ell+1)}{24}$
$=\frac{\ell^{4}(1-2/\ell)(1-1/\ell)(1+1/\ell)}{8}$
$>\frac{\ell^{4}}{16},$
by using $(1-2/\ell)(1-1/\ell)\geq(1-2/12)(1-1/12)>1/2$ for any integer $\ell\geq 12$ and $(1+1/\ell)\geq 1$. Thus, $\lambda$ cannot be admissible.
Then we shall compute $m_{\lambda}(\lambda_{\ell-3})$. We first note
$\mu:=\lambda_{\ell-3}=\lambda-(\alpha_{\ell}+\alpha_{\ell-1}+\alpha_{\ell-2})$
Thus, we have to compute
$m_{\lambda}(\mu)=\frac{(*)}{(\lambda,\lambda)-(\mu,\mu)+2(\lambda-\mu,\delta)},$
where
(*(*)( * ) $=2\sum_{\alpha\in\Phi^{+}}\sum_{i=1}^{\infty}(\mu+i\alpha,\alpha)m_{\lambda}(\mu+i\alpha)$
$=(\mu+(\alpha_{\ell}+\alpha_{\ell-1}+\alpha_{\ell-2}),\alpha_{\ell}+\alpha_{\ell-1}+\alpha_{\ell-2})m_{\lambda}(\mu+(\alpha_{\ell}+\alpha_{\ell-1}+\alpha_{\ell-2}))$
$+(\mu+(\alpha_{\ell}+\alpha_{\ell-1}),\alpha_{\ell}+\alpha_{\ell-1})m_{\lambda}(\mu+(\alpha_{\ell}+\alpha_{\ell-1}))$
$+(\mu+(\alpha_{\ell-1}+\alpha_{\ell-2}),\alpha_{\ell-1}+\alpha_{\ell-2})m_{\lambda}(\mu+(\alpha_{\ell-1}+\alpha_{\ell-2}))$
$+(\mu+\alpha_{\ell},\alpha_{\ell})m_{\lambda}(\mu+\alpha_{\ell})+(\mu+\alpha_{\ell-1},\alpha_{\ell-1})m_{\lambda}(\mu+\alpha_{\ell-1})$
$+(\mu+\alpha_{\ell-2},\alpha_{\ell-2})m_{\lambda}(\mu+\alpha_{\ell-2}).$
By computing -term by term- as in Example 3.3 we may infer that
$m_{\lambda}(\mu+\alpha_{\ell})(\mu+\alpha_{\ell},\alpha_{\ell})$ $=1\cdot 2$
$m_{\lambda}(\mu+\alpha_{\ell-1})(\mu+\alpha_{\ell-1},\alpha_{\ell-1})$ $=1\cdot 2$
$m_{\lambda}(\mu+\alpha_{\ell-2})(\mu+\alpha_{\ell-2},\alpha_{\ell-2})$ $=1\cdot 2$
$m_{\lambda}(\mu+\alpha_{\ell-1}+\alpha_{\ell})(\mu+\alpha_{\ell-1}+\alpha_{\ell},\alpha_{\ell-1}+\alpha_{\ell})$ $=1\cdot 2$
$m_{\lambda}(\mu+\alpha_{\ell-1}+\alpha_{\ell-2})(\mu+\alpha_{\ell-1}+\alpha_{\ell-2},\alpha_{\ell-1}+\alpha_{\ell-2})$ $=1\cdot 2$
$m_{\lambda}(\mu+\alpha_{\ell-2}+\alpha_{\ell-1}+\alpha_{\ell})(\mu+\alpha_{\ell-2}+\alpha_{\ell-1}+\alpha_{\ell},+\alpha_{\ell-2}+\alpha_{\ell-1}+\alpha_{\ell})$ $=1\cdot 2.$
Finally, since
$(\lambda,\lambda)-(\mu,\mu)+2(\lambda-\mu,\delta)=(\lambda,\lambda)-(\mu,\mu)+2(\alpha_{\ell}+\alpha_{\ell-1}+\alpha_{\ell-2},\delta)=6-4+2\cdot 3=8,$
we have:
$m_{\lambda}(\mu)=2\cdot\frac{2\cdot 1+2\cdot 1+2\cdot 1+2\cdot 1+2\cdot 1+2\cdot 1}{8}=2\cdot\frac{12}{8}=3.$
Hence $\lambda=\lambda_{\ell-2}+\lambda_{\ell}$ is not admissible.
Remark 3.5.
It is worth to emphasizing the obvious inequality
$\dim V(\lambda)\geq m_{\lambda}(\mu)l(\mu)\geq l(\mu)$
that we will frequently use for any dominant weight $\mu\prec\lambda$.
Proposition 3.6.
Let $G$ be as above with $\ell\geq 15$. Then, all weights in the following table are admissible.
$\lambda$ dimension of $V(\lambda)$
00 $1$
$\lambda_{1}$, $\lambda_{\ell}$ $\ell+1$
$\lambda_{2}$, $\lambda_{\ell-1}$ $\ell(\ell+1)/2$
$\lambda_{3}$, $\lambda_{\ell-2}$ $(\ell-1)\ell(\ell+1)/6$
$\lambda_{4}$, $\lambda_{\ell-3}$ $(\ell-2)(\ell-1)\ell(\ell+1)/24$
$2\lambda_{1}$, $2\lambda_{\ell}$ $(\ell+2)(\ell+1)/2$
$3\lambda_{1}$, $3\lambda_{\ell}$ $(\ell+3)(\ell+2)(\ell+1)/6$
$\lambda_{1}+\lambda_{\ell}$ $\ell(\ell+2)$
$\lambda_{1}+\lambda_{\ell-1}$, $\lambda_{2}+\lambda_{\ell}$ $(\ell+2)(\ell+1)(\ell-1)/2$
$\lambda_{1}+2\lambda_{\ell}$, $2\lambda_{1}+\lambda_{\ell}$ $(\ell+1)\ell(\ell+3)/2$
$\lambda_{1}+\lambda_{2}$, $\lambda_{\ell-1}+\lambda_{\ell}$ $\ell(\ell+1)(\ell+2)/3$
In addition for $\ell\geq 16$, the weights $4\lambda_{1}$ and $4\lambda_{\ell}$ are also admissible whose associated irreducible representations have dimension $(\ell+4)(\ell+3)(\ell+2)(\ell+1)/24$.
@@ 00 $1$
Firstly, we shall prove that, for $\ell\geq 15$, the following weights
[…]
are admissible. We start remarking that $V(k\lambda_{\ell})$ is irreducible by definition and is contained in the symmetric power tensor product $\operatorname{Sym}^{k}(V)$, where $V$ is the standard irreducible representation of dimension $\ell+1$. Now, let us start by deducing the dimension of the weight space $V(3\lambda_{\ell})$. Note that the dominant weights $\mu\preceq 3\lambda_{\ell}$ are: $3\lambda_{\ell}$, $\lambda_{\ell-1}+\lambda_{\ell}$, and $\lambda_{\ell-2}$. In addition, their stabilizers are contained in the reflection groups of type $A_{\ell-1}$, $A_{\ell-2}$ and $A_{\ell-3}\times A_{2}$, respectively. Since its multiplicities are at least $1$, by formula (2.1) we can observe
$\dim V(3\lambda_{\ell})$ $\geq l(3\lambda_{\ell})+l(\lambda_{\ell-1}+\lambda_{\ell})+l(\lambda_{\ell-2})$
$=(\ell+1)+\ell(\ell+1)+\ell(\ell-1)(\ell+1)/6$
$=(\ell+1)(\ell+2)(\ell+3)/6$
$=\dim\operatorname{Sym}^{3}(V).$
This then implies that $V(3\lambda_{\ell})=\operatorname{Sym}^{3}(V)$ and $\dim V(3\lambda_{\ell})=(\ell+1)(\ell+2)(\ell+3)/6\leq\ell^{4}/6$. By duality we infer the dimensions of $V(3\lambda_{1})$. Similarly, we can deduce that the anti-symmetric power tensor product $\Lambda^{i+1}(V)=V(\lambda_{\ell-i})$ for $i=2,3$ by observing that $V(\lambda_{\ell-i})\subseteq\Lambda^{i+1}(V)$ and
$\dim V(\lambda_{\ell-i})\geq l(\lambda_{\ell-i})=\frac{(\ell+1)!}{(\ell-i)!(i+1)!}=\binom{\ell+1}{i+1}=\dim\Lambda^{i+1}(V).$
[…]
Now, when $\ell\geq 16$ we may check that $\lambda=4\lambda_{\ell}$ is admissible by verifying first that the dominant weights $\mu\preceq 4\lambda_{\ell}$ are: $4\lambda_{\ell}$, $\lambda_{\ell-1}+2\lambda_{\ell}$, $2\lambda_{\ell-1}$, $\lambda_{\ell-2}+\lambda_{\ell-1}$ and $\lambda_{\ell-3}$, and later to check
$\dim V(4\lambda_{\ell})\geq l(4\lambda_{\ell})+l(\lambda_{\ell-1}+2\lambda_{\ell})+l(2\lambda_{\ell-1})+l(\lambda_{\ell-2}+\lambda_{\ell})+l(\lambda_{\ell-3})=\dim\operatorname{Sym}^{4}(V).$
This implies that $V(4\lambda_{\ell})=\operatorname{Sym}^{4}(V)$ and $\dim V(4\lambda_{\ell})=(\ell+1)(\ell+2)(\ell+3)(\ell+4)/24\leq\ell^{4}/16$. Then, by duality, we deduce the dimension of $V(4\lambda_{1})$. Finally, we may note that the remaining weights were studied by Lübeck in [11, §5]. ∎
Remark 3.7.
Notice that the argument used to prove $\Lambda^{i+1}(V)=V(\lambda_{\ell-i})$ for $i=2,3$ (resp. $V(i\lambda_{\ell})=\operatorname{Sym}^{i}(V)$ for $i=3,4$) works for any integer $i$ such that $11\leq i\leq\lfloor\ell/2\rfloor+$ 1. This tell us in particular that $\Lambda^{i+1}(V)$ (resp. $\operatorname{Sym}^{i}(V)$) is irreducible. Moreover, by duality we may observe that
$\dim V(\lambda_{i+1})=\dim\Lambda^{i+1}(V),\quad(\mbox{resp. }\dim V(i\lambda_{1})=\dim\operatorname{Sym}^{i}(V)).$
4. Representations of finite groups of type $A_{\ell}$
In order to determine all admissible weights when $G$ is a finite group of Lie type $A_{\ell}$ over $\mathbb{\overline{F}}_{p}$ we shall proceed by a series of steps to reduce the possibilities according to whether the coefficients are zeroes or not. Let $\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+\cdots+a_{\ell}\lambda_{\ell}$ be a dominant weight. The first step, nullity of coefficients, consists of identifying all indexes $i$ such that $a_{i}=0$. Each of the other reductions will be associated to a nonzero coefficient $a_{i}$, where we determine all the possibilities to the other nonzero coefficients $a_{j}$ reducing us exclusively to admissible weights. Explicitly, we start with.
Nullity of coefficients -- Claim: $a_{i}\neq 0$ only if $i\leq 4$ or $i\geq\ell-3$.
Suppose $a_{i}\neq 0$ for some $i$ satisfying $45\leq i\leq\ell-$ 4. This implies that the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{i-1}\times A_{\ell-i}$. Thus,
$\dim V(\lambda)\geq l(\lambda)=\frac{(\ell+1)!}{i!(\ell+1-i)!}.$
We shall prove that $(\ell+1)!/i!(\ell+1-i)!>\ell^{4}/16$, for any $45\leq i\leq\ell-$ 4. To do this, assume firstly that $i\leq\ell+1-i$. In this case, we need to show that
@@ 4. Representations of finite groups of type $A_{\ell}$
We shall first verify the inequality for $i=5$, that is,
$\frac{(\ell-3)(\ell-2)(\ell-1)\ell(\ell+1)}{5!}>\frac{\ell^{4}}{16}$ $\iff\frac{(\ell-3)(\ell-2)(\ell-1)(\ell+1)}{3\cdot 5}>\frac{\ell^{3}}{2}.$
In fact, since $\ell\geq 15$, we can infer that $(\ell+1)/15>1$ and $151-j/\ell\geq 1-j/$ 15 for $1\leq j\leq 3$. Then, by combining these facts we can deduce
$\frac{(\ell-3)(\ell-2)(\ell-1)(\ell+1)}{3\cdot 5}$ $>(\ell-3)(\ell-2)(\ell-1)$
$\geq\ell^{3}(1-3/\ell)(1-2/\ell)(1-1/\ell)$
$\geq\ell^{3}(1-3/15)(1-2/15)(1-1/15)$
$\geq\ell^{3}\frac{12}{15}\cdot\frac{13}{15}\cdot\frac{14}{15}$
$>\frac{\ell^{3}}{2}.$
Thus, the case $i=5$ follows by multiplying by $\ell/8$.
The general case $i\leq\ell+1-i$ is then deduced from the base case $i=5$ by noting that
@@
$=\left(\frac{(\ell+2-i)(\ell+2-(i+1))\cdots(\ell+2-6)}{i(i-1)\cdots 6}\right)\left(\frac{(\ell-3)(\ell-2)(\ell-1)\ell(\ell+1)}{5!}\right)$
and the first parenthesis is greater than 1 by using the condition $\ell\geq 15$. Finally, the case $\ell+1-i<i\leq\ell$ follows by using a symmetric argument. Therefore, we reduces us to

### v3

@@ 2.3. Dimension and Multiplicity
following [19, §10.2 and §13.3].
We define the multiplicity of $\mu$ in $V$, for any dominant weight $\mu\preceq\lambda$, as
$m_{\lambda}(\mu)=\dim(V_{\mathbb{C}})_{\mu}$
which can be calculated by using the recursive formula [19, §22.3]:
(2.2) $m_{\lambda}(\mu)=\frac{2\sum_{\alpha\in\Phi_{+}}\sum_{i=1}^{\infty}(\mu+i\alpha,\alpha)m_{\lambda}(\mu+i\alpha)}{\parallel\lambda+\delta\parallel^{2}+\parallel\mu+\delta\parallel^{2}}=\frac{2\sum_{\alpha\in\Phi_{+}}\sum_{i=1}^{\infty}(\mu+i\alpha,\alpha)m_{\lambda}(\mu+i\alpha)}{(\lambda,\lambda)-(\mu,\mu)+2(\lambda-\mu,\delta)},$
@@
It follows from [32, (1.10)] (see also [12, Proposition 2.4.3]) that the dimension of $V_{\mathbb{C}}$ is
(2.3) $\dim V_{\mathbb{C}}=\sum_{\mu\in\Pi(\lambda)}m_{\lambda}(\mu)l(\mu),$
Nevertheless, as we mentioned in the Introduction, determining the dimension of $L(\lambda)$ in positive characteristic is much more complicated. In particular, knowing multiplicity $m_{\lambda}(\mu)$ is to compute the rank of the matrix obtained from $C^{(\lambda,\mu)}$ by reduction modulo $p$. Now, given that those calculations are very lengthy even in low rank, computational programs have been a useful tool since the 1970s. In addition, we may observe that the multiplicities obtained by reduction modulo $p$ are not, in general, the same, the appearance of the weights $\mu$ also demands attention. For instance, Premet established that the set of weights for the modules $V$, $L(\lambda)$ is the same if $G$ is of type $B_{\ell}$ and $C_{\ell}$ when $p>3$ and $\lambda$ is $p$-restricted, which are the central ones for us as we will see below.
Thereby, it follows that
(2.4) $\dim L(\lambda)\leq\sum_{\mu\in\Pi(\lambda)}m_{\lambda}(\mu)l(\mu)=\dim V_{\mathbb{C}}.$
We finish this section by specifying the notion of admissible weight mentioned in the Introduction.
Definition 2.3.
Let $G$ be a connected reductive simple algebraic group of rank $\ell$ of type $A,B,C,D$ over $\overline{\mathbb{F}}_{p}$ of simply connected type and $\lambda$ be a dominant weight. We say that $\lambda$ is admissible if
@@
The next three sections will be devoted to determining all the admissible weights for all finite classical groups in their defining characteristic.
3. Representations of finite classical groups of type $A_{\ell}$
In this section we will determine all the admissible weights of type $A_{\ell}$ together with the exact dimension of their associated irreducible representations.
Let $G$ be a finite classical group of type $A_{\ell}$. We shall begin by fixing a set of simple roots of type $A_{\ell}$ and its fundamental weights associated. To this end, let $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{\ell}\}$ be the simple root system of type $A_{\ell}$ given by
@@ 3. Representations of finite classical groups of type $A_{\ell}$
$\lambda_{i}=\frac{1}{\ell+1}[(\ell-i+1)\alpha_{1}+2(\ell-i+1)\alpha_{2}+i(\ell-i+1)\alpha_{i}+i(\ell-i)\alpha_{i+1}+\cdots+i\alpha_{\ell}],$
for $1\leq i\leq\ell$.
We commence by obtaining the following admissible weights.
@@
Let $G$ be as above with $\ell\geq 15$. Then, the weights
$3\lambda,\;4\lambda,\;\lambda_{3},\;\lambda_{4},\;\lambda_{1}+\lambda_{2},\;\lambda_{1}+\lambda_{\ell-1},\;\lambda_{1}+2\lambda_{\ell}$
(and its duals) are admissible.
@@
The weights $3\lambda_{\ell}$ and $4\lambda_{\ell}$ are admissible, since
$\dim L(a\lambda_{\ell})=\binom{\ell+a}{a}<\frac{\ell^{4}}{4},$
for $a=3,4$ and any $p>3$, by appeal to [13, Lemma 2.3]. The admissibility of $\lambda_{3},\;\lambda_{4}$ can be deduced by using item 2 of [13, Lemma 2.3] independently of $p$. On the other hand, item 3, for $2\leq j\leq\ell$ in loc. cit., tells us that
(3.1) $\dim L(\lambda_{1}+\lambda_{j})=j\binom{\ell+2}{j+1}-\epsilon_{p}(j+1)\binom{\ell+1}{j+1}$,
where $\epsilon_{p}(j+1)$ is $1$ if $p$ divides $j+1$ and 0 otherwise. An evident lower bound for the dimension is when $p$ does not divide $j+1$, from which we infer the admissibility of $\lambda_{1}+\lambda_{j}$ for $j=2$ and $j=\ell-1$. Finally, $\lambda_{1}+2\lambda_{\ell}$ is also admissible by appealing to $V(\lambda_{1}+2\lambda_{\ell})$ is contained in $V(\lambda_{1})\otimes V(2\lambda_{\ell})$ whose dimension is clearly less that $\ell^{4}/16$. ∎
Remark 3.2.
An explicit formula for the dimension of $L(\lambda_{1}+2\lambda_{\ell})$
$\dim L(\lambda_{1}+2\lambda_{\ell})=\dim V(\lambda_{1}+2\lambda_{\ell})-\epsilon_{p}(j+2)L(\lambda_{1})=(\ell+1)\left(\frac{\ell(\ell+3)}{2}-\epsilon_{p}(j+2)\right)$
can be found in [12, Lemma 2.4.9].
In contrast with the example above, we observe.
@@
For all $\ell\geq 12$, the weight $\lambda=\lambda_{\ell-2}+\lambda_{\ell}$ is not admissible for any $p$. Concretely, by considering $j=3$ in Eq. (3.1). we may infer that
$\dim L(\lambda_{1}+\lambda_{j})=3\binom{\ell+2}{4}-\binom{\ell+1}{4}=\frac{1}{12}(\ell-1)\ell(\ell+1)(\ell+4)>\frac{\ell^{4}}{16}.$
Now we prove the main result of this section.
Theorem 3.4.
Let $G$ be a finite group of Lie type $A_{\ell}$ over $\mathbb{\overline{F}}_{p}$. If $\ell\geq 15$ then the following list exhausts the admissible $p$-restricted weights.
$\lambda$ dimension of $L(\lambda)$ characteristic
0 11 all
$\lambda_{1}$, $\lambda_{\ell}$ $\ell+1$ all
$\lambda_{2}$, $\lambda_{\ell-1}$ $\ell(\ell+1)/2$ all
$\lambda_{3}$, $\lambda_{\ell-2}$ $(\ell-1)\ell(\ell+1)/6$ all
$\lambda_{4}$, $\lambda_{\ell-3}$ $(\ell-2)(\ell-1)\ell(\ell+1)/24$ all
$2\lambda_{1}$, $2\lambda_{\ell}$ $(\ell+2)(\ell+1)/2$ $p\neq 2$
$3\lambda_{1}$, $3\lambda_{\ell}$ $(\ell+3)(\ell+2)(\ell+1)/6$ $p\neq 2,3$
$\lambda_{1}+\lambda_{\ell}$ $\ell(\ell+2)$ $p\nmid\ell+1$
$\lambda_{1}+\lambda_{\ell}$ $\ell(\ell+2)-1$ $p\mid\ell+1$
$\lambda_{1}+\lambda_{\ell-1}$, $\lambda_{2}+\lambda_{\ell}$ $(\ell+2)(\ell+1)(\ell-1)/2$ $p\nmid\ell$
$\lambda_{1}+\lambda_{\ell-1}$, $\lambda_{2}+\lambda_{\ell}$ $(\ell+2)(\ell+1)(\ell-1)/2-(\ell+1)$ $p\mid\ell$
$\lambda_{1}+2\lambda_{\ell}$, $2\lambda_{1}+\lambda_{\ell}$ $(\ell+1)\ell(\ell+3)/2$ $p\neq 2$,   $p\nmid\ell+2$
$\lambda_{1}+2\lambda_{\ell}$, $2\lambda_{1}+\lambda_{\ell}$ $(\ell+1)\ell(\ell+3)/2-(\ell+1)$ $p\neq 2$,   $p\mid\ell+2$
$\lambda_{1}+\lambda_{2}$, $\lambda_{\ell-1}+\lambda_{\ell}$ $\ell(\ell+1)(\ell+2)/3$ $p\neq 3$
$\lambda_{1}+\lambda_{2}$, $\lambda_{\ell-1}+\lambda_{\ell}$ $\ell(\ell+1)(\ell+2)/3-(\ell-1)\ell(\ell+1)/6$ $p=3$
In addition for $\ell\geq 16$ and $p\neq 2,3$, the weights $4\lambda_{1}$ and $4\lambda_{\ell}$ are also admissible and $p$-restricted, and their associated irreducible representations have dimension $(\ell+4)(\ell+3)(\ell+2)(\ell+1)/24$.
Notice first that all weights in the table above are admissible by appealing to Proposition 3.1 and [24, Proposition 3.2]. We therefore need to prove that the aforementioned list exhausts the admissible weights. To this end, we shall proceed by a series of steps to reduce the possibilities according to whether the coefficients are zeroes or not. Explicitly, the proof begins with.
@@
Let $\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+\cdots+a_{\ell}\lambda_{\ell}$ be a dominant weight. The first step, Nullity of Coefficients, consists on identifying all indexes $i$ such that $a_{i}=0$. Each of the other reductions will be associated to a nonzero coefficient $a_{i}$, where we determine all the possibilities to the other nonzero coefficients $a_{j}$ reducing exclusively to admissible weights. Explicitly, we start with.
Nullity of coefficients - Claim: $a_{i}\neq 0$ only if $i\leq 4$ or $i\geq\ell-3$.
Suppose $a_{i}\neq 0$ for some $i$ that satisfy $5\leq i\leq\ell-4$. This implies that the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{i-1}\times A_{\ell-i}$, by using Remark 2.1. Thus,
$L(\lambda))\geq l(\lambda)=\frac{(\ell+1)!}{i!(\ell+1-i)!}.$
We shall prove that $(\ell+1)!/i!(\ell+1-i)!>\ell^{4}/16$, for any $5\leq i\leq\ell-4$. To do this, assume firstly that $i\leq\ell+1-i$. In this case, we need to show that
@@
We shall first verify the inequality for $i=5$, which is equivalent to
$\frac{(\ell-3)(\ell-2)(\ell-1)(\ell+1)}{3\cdot 5}>\frac{\ell^{3}}{2}.$
Hence, for $\ell\geq 15$, we can observe that
$\frac{(\ell-3)(\ell-2)(\ell-1)(\ell+1)}{3\cdot 5}>(\ell-3)(\ell-2)(\ell-1)\geq\ell^{3}(1-3/15)(1-2/15)(1-1/15)>\frac{\ell^{3}}{2}$
by appealing to $(1-3/\ell)(1-2/\ell)(1-1/\ell)>1/2$, since $1-j/\ell\geq 1-j/15$ for $1\leq j\leq 3$. Thus, the case $i=5$ follows by multiplying by $\ell/8$. The general case $i\leq\ell+1-i$ is then deduced from the base case $i=5$ by noting that
@@
$\left(\frac{(\ell+2-i)(\ell+2-(i+1))\cdots(\ell+2-6)}{i(i-1)\cdots 6}\right)\left(\frac{(\ell-3)(\ell-2)(\ell-1)\ell(\ell+1)}{5!}\right)$
and the first parenthesis is greater than 1 by using the condition $\ell\geq 15$. Finally, the case $\ell+1-i<i\leq\ell$ follows by using a symmetric argument. Therefore, we reduce to